A short-run average total cost or average cost curve is given by the
equation"
SAC (Q)= 12/Q+0.14Q

Derive expressions for the corresponding short-run total cost, short-run average variable
cost, and short-run average fixed cost.
1
Expert's answer
2020-03-31T08:58:13-0400

If SAC (Q) = 12/Q + 0.14Q, then:

short-run average fixed cost is "SAFC = 12\/Q,"

short-run average variable cost is "SAVC = 0.14Q,"

short-run total cost is "STC = SAC*Q = 12 + 0.14Q^2."